OpenStudy (anonymous):

Solve the initial value problem \(\large y ' =\dfrac{x^2-y^2 }{xy}\) with y(2)=3 .

ganeshie8 (ganeshie8):

Note that the expression on right hand side is of form \(\large f(kx, ky) = k^nf(x, y) \)

ganeshie8 (ganeshie8):

try the substitution \(\large \rm y = vx\)

OpenStudy (anonymous):

i did but i got the wrong answer

ganeshie8 (ganeshie8):

\(\large y = vx \implies y' = v'x + v\)

ganeshie8 (ganeshie8):

the equation becomes \(\large v'x + v = \dfrac{x^2-(vx)^2}{x(vx)}\) \(\large v'x + v = \dfrac{1-v^2}{v}\)

ganeshie8 (ganeshie8):

yes ?

OpenStudy (anonymous):

i think when i put the value of y(2)=3 , i made mistake

ganeshie8 (ganeshie8):

whats teh general solution you got ?

OpenStudy (anonymous):

1x (x 4 +562 ) 12

OpenStudy (anonymous):

1/x(X^(4)+56)/2)^(1/2)

ganeshie8 (ganeshie8):

is there some change in the question ? whats the original equation ?

ganeshie8 (ganeshie8):

\(\large y ' =\dfrac{x^2-y^2 }{xy} \) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

with y(2)= 3

ganeshie8 (ganeshie8):

Alright, lets see if i get the same solution as yours : \(\large \dfrac{dv}{dx}x + v = \dfrac{1-v^2}{v}\) \(\large \dfrac{dv}{dx}x = \dfrac{1-2v^2}{v} \) \(\large \int\dfrac{v}{1-2v^2}dv = \int\dfrac{1}{x}dx \) \(\large -\frac{1}{4}\ln|1-2v^2| = \ln|cx| \) \(\large (1-2v^2)^{-1/4} = cx\) \(\large 1-2v^2 = Cx^{-4}\) \(\large v^2 =\dfrac{1- Cx^{-4}}{2}\) \(\large v^2 =\dfrac{x^4 - C}{2x^4}\)

ganeshie8 (ganeshie8):

plugging in v = y/x back, you get : \(\large (y/x)^2 = \dfrac{x^4-C}{2x^4}\) \(\large y^2 = \dfrac{x^4-C}{2x^2}\)

ganeshie8 (ganeshie8):

plugin the initial value x = 2, y=3 and solve C

OpenStudy (anonymous):

i didnt get why you put v^2 as (x/y)^2 (y/x) 2 =x 4 −C2x 4

OpenStudy (anonymous):

second last step

OpenStudy (anonymous):

2y^(2)X^(2)=x^(4)-c c=-x^(2)(2y^2-x^2)

OpenStudy (anonymous):

so i got c=-56

OpenStudy (anonymous):

so my answer will be y=Square root(x^(4)+56)/(2x^(4))

OpenStudy (anonymous):

Determine which of the following pairs of functions are linearly independent. f(t)=e ^(λt) cos(μt), and g(t)=e ^(λt) sin(μt),μ≠0

ganeshie8 (ganeshie8):

that looks good!

ganeshie8 (ganeshie8):

for posting a new question, you need to close this question and open a new thread